# Elliptic asymptotic representation of the fifth Painlevé transcendents (corrected version)

### v8

@@ A Stokes curve is defined by
where $\lambda_{*}$ is a turning point [11]. This curve connects $\lambda_{*}$ to another turning point, $\pm e^{i\phi}$ or $\infty.$
In what follows, the Stokes graph is discussed under the supposition $a_{\phi}(t)\to A_{\phi}$ as $t\to\infty$ (cf. (4.3)), where $A_{\phi}$ is a unique solution of the Boutroux equations (2.3). Suppose that $0<|\phi|<\pi/2$. Let us consider the limit Stokes graph for $t=\infty$, in which the limit turning points $\lambda_{1,\,2}(\infty)$ are also denoted by the same symbols $\lambda_{1,\,2}$ as for $t\not$$\infty$. Set $\mathbb{P}_{+}^{\infty}\cup\mathbb{P}_{-}^{\infty}=\lim_{t\to\infty}\mathbb{P}_{+}\cup\mathbb{P}_{-}$, which is a two-sheeted Riemann surface glued along the cuts $e^{i\phi},\lambda_{2}], [$$\lambda_{1},e^{i\phi}]$. As long as the turning points do not coalesce, the limit Stokes curve for $t=\infty$ reflects the Boutroux equations. Then this limit Stokes graph on $\mathbb{P}_{+}^{\infty}$ is considered to be as in Figure 3.2 (b), (c) (cf. Proposition 8.17), in which, if $0<\phi<\pi/2,$ the Stokes curves connect $\lambda_{1}$ to $e^{i\phi}$, $\lambda_{2}$ and $i\infty,$ and $\lambda_{2}$ to $e^{i\phi}$ and $i\infty.$
@@ 3.4. WKB-solution
To remove $R=T^{-1}T_{\lambda}-\mathrm{diag}T^{-1}T_{\lambda}$ set
@@ 3.5. Local solutions around turning points
$T_{\iota}^{-1}\mathcal{B}_{1}(t,\lambda)T_{\iota}=\begin{pmatrix}ic_{3}^{-1}(c_{2}\delta_{1}-c_{1}\delta_{2})&(c_{1}+ic_{2})^{-1}(c_{1}\delta_{1}+c_{2}\delta_{2}-c_{3}\delta_{3})\\ (c_{1}-ic_{2})^{-1}(c_{1}\delta_{1}+c_{2}\delta_{2}+c_{3}\delta_{3})&-ic_{3}^{-1}(c_{2}\delta_{1}-c_{1}\delta_{2}),\end{pmatrix}$
@@ 4.1. Stokes graph
When $t\to\infty,$ the turning points $\lambda^{0}_{1,2}$ tend to $\pm e^{i\phi}$. As in Figure 4.1 (a), (b), symbols are assigned to the Stokes curves of the limit Stokes graph on $\mathbb{P}_{+}^{\infty}$. For $\pi/2<\phi<0,$ let $\mathbf{c}^{\infty}_{2},$ $\hat{\mathbf{c}}^{\infty}_{1},$ $\mathbf{c}_{0}$ be Stokes curves on $\mathbb{P}_{+}^{\infty}$ connecting $\lambda_{2}$ to $i\infty,$ $\lambda_{1}$ to $i\infty,$ $\lambda_{2}$ to $\lambda_{1}$, respectively, and, for $0<\phi<\pi/2$, $\mathbf{c}^{\infty}_{1},$ $\hat{\mathbf{c}}^{\infty}_{2},$ $\mathbf{c}_{0}$ connecting $\lambda_{1}$ to $i\infty,$ $\lambda_{2}$ to $i\infty$, $\lambda_{2}$ to $\lambda_{1}$, respectively.
@@ 4.2. Connection matrices
In the domain $\mathcal{D}^{0}_{+}=\mathbb{C}\setminus[-\infty,-e^{i\phi}]$ with $|\arg\lambda|<\pi$, considering the projection of the Stokes curve $\gamma_{1}=(-\mathbf{c}^{\infty}_{1})\cup(-\mathbf{c}_{0})\cup\hat{\mathbf{c}}^{\infty}_{2}$ issuing from $e^{\pi i/2}\infty$ and ending in $e^{-\pi i/2}\infty,$ we analogously obtain
@@
where $\Gamma^{\infty}_{\infty\,1}$ is a connection matrix such that $Y^{\gamma_{1}}\Gamma^{\infty}_{\infty\,1}=Y_{1}$ in $\mathcal{D}_{+}(-\pi/2):=\{|\arg\lambda+\pi/2|<\pi/4,\,\,|\lambda|>2025\}\subset\mathcal{D}^{0}_{+}$. The connection matrices $\Gamma^{\infty}_{\infty\,1}$ and $\Gamma^{\infty}_{\infty\,2}$ are constructed by combining matching procedures.
Let us explain the calculation of $\Gamma^{\infty}_{\infty 2}$ in the case $0<\phi<\pi/2.$ To discuss according to the justification scheme of Kitaev [26], suppose that $a_{\phi}=a_{\phi}(t)$ is given by (3.10) with $(y,y^{*})=(y(t),y^{*}(t))$ not necessarily solving (PV), and that
@@
for $t\in S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$ with given $\kappa_{0},$ given small $\delta_{1}$ and sufficiently large $t^{\prime}_{\infty}$. Here $A_{\phi}$ is a unique solution of the Boutroux equation (2.3), and
$S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})=\{t\,;\,\,\mathrm{Re\,}t>t^{\prime}_{\infty},|\mathrm{Im\,}t|<\kappa_{0},|y^{*}(t)|+|y(t)|+|y(t)|^{-1}+|y(t)-1|^{-1}<\delta_{1}^{-1}\}.$
@@
In Propositions 3.3 and 3.4, set $\delta=\delta^{\prime}=1/4-\varepsilon$, $0<\varepsilon<1/4$ being arbitrary. Then both propositions apply to the annulus
@@
The annulus $|t|^{-14/27-(2/3)\varepsilon^{\prime}}\ll|\lambda-\lambda_{\iota}|\ll|t|^{-14/27+(1/3)\varepsilon^{\prime}}$ of the early version, which is contained in $\mathcal{A}_{\varepsilon}$ with $\varepsilon=1/36+\varepsilon^{\prime}$ for $\delta=\delta^{\prime}=2/9-\varepsilon^{\prime}=1/4-\varepsilon$, is also available in our calculation below, though the derivation is based on an inaccurate argument.
In what follows we set
$\delta=1/4-\varepsilon.$
@@
(b) For $\Psi_{\infty}(\lambda)$ and $\Phi_{1}(\lambda)$ (cf. Proposition 3.4) in the annulus $\mathcal{A}_{\varepsilon}$ set $\Psi_{\infty}(\lambda)=\Phi_{1}(\lambda)\Gamma_{\mathrm{b}}$ along $\mathbf{c}^{\infty}_{1}.$ We may suppose that the curve $(2\kappa)^{1/3}(\lambda-\tilde{\lambda}_{1})=t^{-2/3}(\zeta+O(t^{-1/3}))$ with $\lambda\in\mathbf{c}^{\infty}_{1}$ enters the sector $\Sigma_{1}:$ $|\arg\zeta-\pi/3|<2\pi/3,$ and that $\Sigma_{1}$ does not intersect the cut $e^{i\phi},\lambda_{1}].$ Write $K^{-1}=2(2\kappa)^{-1/3}(d_{1}-id_{2}).$ Then, by Propositions 3.3 and 3.4,
@@
$\Gamma_{\mathrm{f}}=({\zeta}^{\prime}_{2})^{1/4}(I+O(t^{-\delta}))\tilde{C}^{\prime}_{3}(\tilde{\lambda}^{\prime}_{2})^{-1}\begin{pmatrix}1&0\\ 0&-c_{0}\end{pmatrix}$
with ${\zeta}^{\prime}_{2}\asymp{\lambda}^{\prime}_{2}-\lambda_{2}.$
@@
The matrix $\Gamma^{\infty}_{\infty 1}$ is calculated by using the same Stokes graph of Figure 4.1 (b) on the plane $\mathbb{C}\setminus[-\infty,-e^{i\phi}]$ containing the sector $\pi<\arg\lambda<\pi,$ $|\lambda|>2025.$ Note that, in this case, the curve $\hat{\mathbf{c}}^{\infty}_{2}$ tends to $e^{-\pi i/2}\infty$, and denote it by $\hat{\mathbf{c}}^{\infty}_{2*}.$ The calculation of $\Gamma^{\infty}_{\infty 1}$ begins with setting $Y_{1}(\lambda)=\hat{\Psi}^{*}_{\infty}(\lambda)\Gamma^{*}_{\mathrm{a}},$ where $\hat{\Psi}^{*}_{\infty}(\lambda)$ is the WKB solution along $\hat{\mathbf{c}}^{\infty}_{2*}$ tending to $e^{-\pi i/2}\infty$. Repeating step by step matchings along $\gamma_{1}=(-\mathbf{c}_{1}^{\infty})\cup\mathbf{c}_{0}\cup\hat{\mathbf{c}}^{\infty}_{2*}$ we have
@@
$\hat{J}_{1}=\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\in\hat{\mathbf{c}}^{\infty}_{1}\end{subarray}}\Bigl(\int^{\lambda}_{\lambda_{1}}\Lambda_{3}(\tau)d\tau-\frac{1}{4}(t\lambda-2\theta_{\infty}\log\lambda)\sigma_{3}\Bigr)$
@@ 5. Asymptotics of monodromy data
$w=w(z)=\sqrt{(1-z^{2})(a_{\phi}-z^{2})}$
and $\tilde{g}_{2}(t,z)\ll 1$ if $|z^{2}-1|^{-1},$ $|z^{2}-a_{\phi}|^{-1}\ll 1.$ Let $z(\mathbb{P})=z(\mathbb{P}_{+})\cup z(\mathbb{P}_{-})$ be the image of $\mathbb{P}=\mathbb{P}_{+}\cup\mathbb{P}_{-}$ under the map (5.1), and set $(\Pi,\Pi_{+},\Pi_{-})=(z(\mathbb{P}),z(\mathbb{P}_{+}),z(\mathbb{P}_{-}))|_{*}$, where the modification $|_{*}$ denotes the replacement of the cuts $z_{1},z^{0}_{1}]\mapsto[a^{1/2}_{\phi},1], [$$z^{0}_{2},z_{2}]\mapsto[-1,-a^{1/2}_{\phi}].$ Then $\Pi$ is the elliptic curve defined by $w(z)$, which consists of two copies of $z$-plane glued along the cuts $1,-a_{\phi}^{1/2}]$ and $a_{\phi}^{1/2},1]$. In (5.2) each square root is such that $a^{-1/2}_{\phi}\sqrt{(a_{\phi}-z^{2})/(1-z^{2})}$, $a^{-1/2}_{\phi}w(z)\to\pm 1$ as $z\to 0$ on $\Pi_{\pm}.$ The characteristic root $\mu=\mu(z)$ itself is an algebraic function on $z(\mathbb{P})$ such that $a^{-1/2}_{\phi}\mu(z)\to\pm 1/4$ as $z\to 0$ on $z(\mathbb{P}_{\pm}).$
Let $\mathbf{a}$ and $\mathbf{b}$ be the cycles on $\Pi$ descried as in Figure 5.1.
$a_{\phi}^{1/2}11$$_{\phi}^{1/2}𝐚\mathbf{a}$$\mathbf{b}$$\Pi_{+}$
Figure 5.1. Cycles $\mathbf{a},$ $\mathbf{b}$ on $\Pi$
We remark that $\mathbf{a}$ and $\mathbf{b}$ may also be regarded as those on $z(\mathbb{P})$, if $t$ is sufficiently large and if the distance between $\mathbf{a}\cup\mathbf{b}$ and $\{\pm 1\}\cup\{\pm a_{\phi}^{1/2}\}$ is $\gg 1$. We use the same symbols $\mathbf{a}$ and $\mathbf{b}$ as in Figure 2.2, provided that $A^{1/2}_{\phi}=\lim_{t\to\infty}a^{1/2}_{\phi}(t),$ which will not cause confusions.

### v9

@@ A Stokes curve is defined by
where $\lambda_{*}$ is a turning point [11]. This curve connects $\lambda_{*}$ to another turning point, $\pm e^{i\phi}$ or $\infty.$ The Stokes graph consists of Stokes curves, turning points and singular points.
Our WKB analysis will be carried out under the supposition $a_{\phi}(t)\to A_{\phi}$ as $t\to\infty$ (cf. (4.3)), where $A_{\phi}$ is a unique solution of the Boutroux equations (2.3). Suppose that $0<|\phi|<\pi/2$. Let us consider the limit Stokes graph for $t=\infty$, in which the limit turning points $\lambda_{1,\,2}(\infty)$ are also denoted by the same symbols $\lambda_{1,\,2}$ as for $t\not$$\infty$. Set $\mathbb{P}_{+}^{\infty}\cup\mathbb{P}_{-}^{\infty}=\lim_{t\to\infty}\mathbb{P}_{+}\cup\mathbb{P}_{-}$, which is a two-sheeted Riemann surface glued along the cuts $e^{i\phi},\lambda_{2}], [$$\lambda_{1},e^{i\phi}]$ with $\lambda_{1,\,2}=\lambda_{1,\,2}(\infty)=\pm e^{i\phi}A_{\phi}^{1/2}$ and $\lambda^{0}_{1,\,2}(\infty)=\pm e^{i\phi}$. As long as the turning points do not coalesce, the limit Stokes curve for $t=\infty$ reflects the Boutroux equations. Then this limit Stokes graph on $\mathbb{P}_{+}^{\infty}$ is considered to be as in Figure 3.2 (b), (c) (cf. Proposition 8.17), in which, if $0<\phi<\pi/2,$ the limit Stokes curves connect $\lambda_{1}$ to $e^{i\phi}$, $\lambda_{2}$ and $i\infty,$ and $\lambda_{2}$ to $e^{i\phi}$ $\lambda_{1}$ and $i\infty.$
@@ 3.4. WKB-solution
To remove the off-diagonal part $R=T^{-1}T_{\lambda}-\mathrm{diag}T^{-1}T_{\lambda}$ set
@@ 3.5. Local solutions around turning points
$T_{\iota}^{-1}\mathcal{B}_{1}(t,\lambda)T_{\iota}=\begin{pmatrix}ic_{3}^{-1}(c_{2}\delta_{1}-c_{1}\delta_{2})&(c_{1}+ic_{2})^{-1}(c_{1}\delta_{1}+c_{2}\delta_{2}-c_{3}\delta_{3})\\ (c_{1}-ic_{2})^{-1}(c_{1}\delta_{1}+c_{2}\delta_{2}+c_{3}\delta_{3})&-ic_{3}^{-1}(c_{2}\delta_{1}-c_{1}\delta_{2})\end{pmatrix}$,
@@ 4.1. Stokes graph
Recall the Riemann surface $\mathbb{P}_{+}^{\infty}\cup\mathbb{P}_{-}^{\infty}$ glued along the cuts $e^{i\phi},\lambda_{2}]$ and $\lambda_{1},e^{i\phi}]$ with $\lambda_{2,\,1}=\mp e^{i\phi}A_{\phi}^{1/2}$ for $t=\infty.$ As in Figure 4.1 (a), (b), symbols are assigned to the Stokes curves of the limit Stokes graph on $\mathbb{P}_{+}^{\infty}$. For $\pi/2<\phi<0,$ let $\mathbf{c}^{\infty}_{2},$ $\hat{\mathbf{c}}^{\infty}_{1},$ $\mathbf{c}_{0}$ be Stokes curves on $\mathbb{P}_{+}^{\infty}$ connecting $\lambda_{2}$ to $i\infty,$ $\lambda_{1}$ to $i\infty,$ $\lambda_{2}$ to $\lambda_{1}$, respectively, and, for $0<\phi<\pi/2$, $\mathbf{c}^{\infty}_{1},$ $\hat{\mathbf{c}}^{\infty}_{2},$ $\mathbf{c}_{0}$ connecting $\lambda_{1}$ to $i\infty,$ $\lambda_{2}$ to $i\infty$, $\lambda_{2}$ to $\lambda_{1}$, respectively.
@@ 4.2. Connection matrices
In the domain $\mathcal{D}^{0}_{+}=\mathbb{C}\setminus[-\infty,-e^{i\phi}]$ with $|\arg\lambda|<\pi$, for the Stokes curve $\gamma_{1}=(-\mathbf{c}^{\infty}_{1})\cup(-\mathbf{c}_{0})\cup\hat{\mathbf{c}}^{\infty}_{2}$ on $\mathbb{P}_{+}^{\infty}$, considering the projection to $\mathcal{D}^{0}_{+}$ issuing from $e^{\pi i/2}\infty$ and ending in $e^{-\pi i/2}\infty,$ we analogously obtain
@@
where $\Gamma^{\infty}_{\infty\,1}$ is a connection matrix such that $Y^{\gamma_{1}}\Gamma^{\infty}_{\infty\,1}=Y_{1}$ in $\mathcal{D}_{+}(-\pi/2):=\{|\arg\lambda+\pi/2|<\pi/4,\,\,|\lambda|>2025\}\subset\mathcal{D}^{0}_{+}$. Our concern is finding the connection matrices $\Gamma^{\infty}_{\infty\,1}$ and $\Gamma^{\infty}_{\infty\,2}$.
Let us calculate $\Gamma^{\infty}_{\infty 2}$ in the case $0<\phi<\pi/2.$ To discuss according to the justification scheme of Kitaev [26], suppose that $a_{\phi}=a_{\phi}(t)$ is given by (3.10) with $(y,y^{*})=(y(t),y^{*}(t))$ not necessarily solving (PV), and that
@@
for $t\in S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$ with given $\kappa_{0},$ given small $\delta_{1}$ and sufficiently large $t^{\prime}_{\infty}$. Here $A_{\phi}$ is a unique solution of the Boutroux equations (2.3), and
$S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})=\{t\,|\,\,\mathrm{Re\,}t>t^{\prime}_{\infty},|\mathrm{Im\,}t|<\kappa_{0},|y^{*}(t)|+|y(t)|+|y(t)|^{-1}+|y(t)-1|^{-1}<\delta_{1}^{-1}\}.$
@@
In Propositions 3.4 and 3.5, set $\delta=\delta^{\prime}=1/4-\varepsilon$, $0<\varepsilon<1/4$ being arbitrary. Then both propositions apply to the annulus
@@
The annulus $|t|^{-14/27-(2/3)\varepsilon^{\prime}}\ll|\lambda-\lambda_{\iota}|\ll|t|^{-14/27+(1/3)\varepsilon^{\prime}}$ of the early version, which is contained in $\mathcal{A}_{\varepsilon}$ with $\varepsilon=1/36+\varepsilon^{\prime}$ for $\delta=\delta^{\prime}=2/9-\varepsilon^{\prime}=1/4-\varepsilon$, is also available in our calculation below, though the derivation of $\frac{14}{27}-\frac{2}{3}\varepsilon^{\prime}$, $\frac{14}{27}+\frac{1}{3}\varepsilon^{\prime}$ is based on an inaccurate argument.
In what follows we set $\delta=1/4-\varepsilon.$
@@
(b) For $\Psi_{\infty}(\lambda)$ and $\Phi_{1}(\lambda)$ (cf. Proposition 3.5) in the annulus $\mathcal{A}_{\varepsilon}$ set $\Psi_{\infty}(\lambda)=\Phi_{1}(\lambda)\Gamma_{\mathrm{b}}$ along $\mathbf{c}^{\infty}_{1}.$ We may suppose that the curve $(2\kappa)^{1/3}(\lambda-\tilde{\lambda}_{1})=t^{-2/3}(\zeta+O(t^{-1/3}))$ with $\lambda\in\mathbf{c}^{\infty}_{1}$ enters the sector $\Sigma_{1}:$ $|\arg\zeta-\pi/3|<2\pi/3,$ and that $\Sigma_{1}$ does not intersect the cut $e^{i\phi},\lambda_{1}].$ Write $K^{-1}=2(2\kappa)^{-1/3}(d_{1}-id_{2}),$ where $d_{k}=b_{k}(\lambda_{1})$ as in (2). Then, by Propositions 3.4 and 3.5,
@@
$\Gamma_{\mathrm{f}}=({\zeta}^{\prime}_{2})^{1/4}(I+O(t^{-\delta}))\tilde{C}^{\prime}_{3}(\tilde{\lambda}^{\prime}_{2})^{-1}\begin{pmatrix}1&0\\ 0&-c_{0}\end{pmatrix}$,
where ${\zeta}^{\prime}_{2}\asymp{\lambda}^{\prime}_{2}-\lambda_{2}$ and $c_{k}=b_{k}(\lambda_{2}).$
@@
The matrix $\Gamma^{\infty}_{\infty 1}$ is calculated by using the same Stokes graph of Figure 4.1 (b) on the plane $\mathcal{D}_{+}^{0}=\mathbb{C}\setminus[-\infty,-e^{i\phi}]$ containing the sector $\pi<\arg\lambda<\pi,$ $|\lambda|>2025.$ Note that, in this case, the curve $\hat{\mathbf{c}}^{\infty}_{2}\subset\mathcal{D}_{+}^{0}$ tends to $e^{-\pi i/2}\infty$, and denote it by $\hat{\mathbf{c}}^{\infty}_{2*}.$ The calculation of $\Gamma^{\infty}_{\infty 1}$ begins with setting $Y_{1}(\lambda)=\hat{\Psi}^{*}_{\infty}(\lambda)\Gamma^{*}_{\mathrm{a}},$ where $\hat{\Psi}^{*}_{\infty}(\lambda)$ is the WKB solution along $\hat{\mathbf{c}}^{\infty}_{2*}$ tending to $e^{-\pi i/2}\infty$. Note that $\hat{\Psi}^{*}_{\infty}(\lambda)$ has the same asymptotic form as of $\hat{\Psi}_{\infty}(\lambda)$. Repeating step by step matchings along $\gamma_{1}=(-\mathbf{c}_{1}^{\infty})\cup\mathbf{c}_{0}\cup\hat{\mathbf{c}}^{\infty}_{2*}$, we have
@@
$\hat{J}_{1}=\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\in\hat{\mathbf{c}}^{\infty}_{1}\end{subarray}}\Bigl(\int^{\lambda}_{\lambda_{1}}\Lambda_{3}(\tau)d\tau-\frac{1}{4}(t\lambda-2\theta_{\infty}\log\lambda)\sigma_{3}\Bigr),$
@@ 5. Asymptotics of monodromy data
$w=w(z)=w(a_{\phi},z)=\sqrt{(1-z^{2})(a_{\phi}-z^{2})}$
and $\tilde{g}_{2}(t,z)\ll 1$ if $|z^{2}-1|^{-1}+|z^{2}-a_{\phi}|^{-1}\ll 1.$ The image of $\mathbb{P}=\mathbb{P}_{+}\cup\mathbb{P}_{-}$ under the map (5.1) is given by $\Pi^{*}=\Pi^{*}_{+}\cup\Pi^{*}_{-}$ with $\Pi_{\pm}^{*}=z(\mathbb{P}_{\pm})$ glued along the cuts $z_{2}^{0}(t),z_{2}(t)]$ and $z_{1}(t),z_{1}^{0}(t)]$, $\mathbb{P}$ being defined in Subsection 3.3. Then $\mu=\mu(z)$ is an algebraic function on $\Pi^{*}.$ The elliptic curve defined by $w(z)$ is given by the two-sheeted Riemann surface $\Pi_{a_{\phi}}=\Pi_{a_{\phi},+}\cup\Pi_{a_{\phi},-}$ glued along the cuts $1,-a_{\phi}^{1/2}]$, $a_{\phi}^{1/2},1]$ satisfying $z_{2,\,1}(t)=\mp a_{\phi}^{1/2}+O(t^{-1})$, $z_{2,\,1}^{0}(t)=\mp 1+O(t^{-2}).$ Each square root in (5.2) is such that $\sqrt{(a_{\phi}-z^{2})/(1-z^{2})}\to 1,$ $z^{-2}w(z)\to-1$ as $z\to\infty$ on $\Pi_{a_{\phi},+}$, and then $a_{\phi}^{-1/2}\sqrt{(a_{\phi}-z^{2})/(1-z^{2})}$,   $a_{\phi}^{-1/2}w(z)\to 1$ as $z\to 0$ on $\Pi_{a_{\phi},+}$. The elliptic curve $\Pi_{A_{\phi}}=\Pi_{+}\cup\Pi_{-}$ of Section 2 is the image of $\mathbb{P}_{+}^{\infty}\cup\mathbb{P}_{-}^{\infty}=\lim_{t\to\infty}\mathbb{P}_{+}\cup\mathbb{P}_{-}$ under the map (5.1).
$a_{\phi}^{1/2}11$$_{\phi}^{1/2}𝐚\mathbf{a}$$\mathbf{b}$$\Pi_{a_{\phi},+}$
Figure 5.1. Cycles $\mathbf{a},$ $\mathbf{b}$ on $\Pi_{a_{\phi}}$
Recall the cycles $\mathbf{a}$ and $\mathbf{b}$ on $\Pi_{A_{\phi}}$ descried in Figure 2.2. We remark that $\mathbf{a}$ and $\mathbf{b}$ may also be regarded as those on $\Pi_{a_{\phi}}$ as in Figure 5.1, if $t$ is sufficiently large and if the distance between $\mathbf{a}\cup\mathbf{b}$ and $\{\pm 1\}\cup\{\pm a_{\phi}^{1/2}\}$ is $\gg 1$. We use the same symbols $\mathbf{a}$ and $\mathbf{b}$ as in Figure 2.2, provided that $A^{1/2}_{\phi}=\lim_{t\to\infty}a^{1/2}_{\phi}(t),$ which will not cause confusions.